OpenStudy (anonymous):

Jack D. Ripper flipped out after missing a Must-Do-It question for the third time on his Minds On Physics assignment. Outraged by the futility of his efforts, he flings a 4.0-gram pencil across the room. The pencil lodges into a 221.0-gram Sponge Bob doll which is at rest on a countertop. Once in motion, the pencil/doll combination slide a distance of 11.9 cm across the countertop before stopping. The coefficient of friction between the doll and the countertop is 0.325. Determine the speed at which the pencil is moving prior to striking Sponge Bob.

OpenStudy (turingtest):

the two equations to use here are\[F=ma\]and\[v_f^2=v_0^2+2ad\](for some reason people always forget that kinematic equation I wrote, but it's so useful)

OpenStudy (anonymous):

but would a be 9.8m/s^2? It travels horizontally, so I thought gravity didn't matter...

OpenStudy (turingtest):

first, how are we going to represent the first event: the transfer of momentum from the pencil to the pencil-spongebob comination?

OpenStudy (anonymous):

p=mv... p=(0.225kg)(v)

OpenStudy (turingtest):

ok I took the srong approach this is an energy problem...

OpenStudy (turingtest):

wrong*

OpenStudy (anonymous):

no it's a momentum problem. http://www.physicsclassroom.com/calcpad/momentum/problems.cfm number 28

OpenStudy (turingtest):

really? because this is what happened to me when I tried it as a momentum problem\[Ft=J=\Delta p\]the initial momentum of the object is\[p_0=(m_1+m_2)v_0\]and the only force acting is friction\[f=\mu N=\mu (m_1+m_2)g\]since the object comes to a rest we have that\[p_0-ft=0\]but since we don't have the time we need the fact that the force of friction acting over that distance will remove all the kinetic energy from the object, so we need to use energy it seems

OpenStudy (turingtest):

so in term of energy that would be\[\frac12(m_1+m_2)v_0^2-fd=0\]and now we can solve for vo, from which we can go back to momentum and get v0 for the pencil from before the collision

OpenStudy (turingtest):

so we use a combination of energy and momentum I think here

OpenStudy (anonymous):

I'm not sure how u came to the conclusion that initial momentum minus Ft=0. But i'm with you so far...and i'm not sure, maybe we could use energy.

OpenStudy (turingtest):

it's impulse I'm not saying the initial momentum is zero (though I could since I can change reference frames at any time) I'm just saying the force times time is change in momentum in this case the object ends at rest, so when the impulse is applied the momentum goes to zero but as I said, we don't know the time, so go the other way I wrote about

OpenStudy (turingtest):

let's just take it from the sponge-pencil moment of combination... what is the kinetic energy of that system?

OpenStudy (anonymous):

0J

OpenStudy (turingtest):

no, the object is moving we won't have a number as an answer (really you should wait and put numbers in at the end when possible)

OpenStudy (anonymous):

we don't know speed though. so KE=1/2mv^2

OpenStudy (turingtest):

where m is really two masses and what is the force of friction?

OpenStudy (anonymous):

i got 0.717 for the force of friction.

OpenStudy (turingtest):

in symbols?

OpenStudy (turingtest):

let's try to save all numbers until the very end if we can

OpenStudy (anonymous):

Ff=μN Ff=0.325*2.205 Ff=0.717

OpenStudy (turingtest):

how about just leaving it as f=μN=μmg for now? it looks much prettier and is easier to see what we are talking about and identify the parts

OpenStudy (anonymous):

lol ok

OpenStudy (turingtest):

also, it is not uncommon for things to cancel algebraically, so unless you want to keep doing the same redundant calculations in you calculator again and again (times mass, divided by mass, then times mass again for example) I advise you to do it the cool way now back to the matter at hand, so what do we need to make this solveable? we need to think about the definition of work, which is what?

OpenStudy (anonymous):

W=Fd

OpenStudy (turingtest):

yes, that is one and we know the force acting on the doll, and the distance over which it acts and what has this got to do with the kinetic energy of the object?

OpenStudy (anonymous):

W=0.717(11.9) W=8.5J...Doesn't that mean that KE=8.5J? But this is the work that friction does on the object. correct?

OpenStudy (turingtest):

ah numbers! I don't want to reach for my calculator... oh wait, I do read you say "Doesn't that mean that KE=... But this is the work that friction does on the object. correct?" the answer to that is a big "yes" the other definition for work besides the one you gave a simple version of is that work equal the change in energy of an object over a distance

OpenStudy (turingtest):

so that is the important part, that\[W=fd=\Delta K\]in general\[W=\Delta K+\Delta U\]but here there is no change in potential energy, so \(\Delta U=0\)

OpenStudy (turingtest):

so now you should be able to deduce the initial velocity of the sponge-pencil object

OpenStudy (anonymous):

So how does that give us kinetic energy? Because we still don't know v... Do I substitute Fd=mv^2? I'm confused as to how to proceed...

OpenStudy (turingtest):

you know m, d, and F, solve for v

OpenStudy (anonymous):

0.085J=mv^2? but it's change in v right

OpenStudy (turingtest):

well yes, that is a good observation, that is not necessarily really \(v_0\) but is actually \(\Delta v\) however in this case what is the final velocity?

OpenStudy (anonymous):

0. but that would give us 0.085=0

OpenStudy (anonymous):

oh no, no it wouldn't. sry.

OpenStudy (anonymous):

0.085=0.225(v^2) vo=0.615m/s

OpenStudy (turingtest):

ok now I will check you numbers, let me get a calculator in the meantime, you realize what happened exactly here?

OpenStudy (anonymous):

Yup. I got it, thanks so far

OpenStudy (turingtest):

we have that\[fd=\Delta K=\frac12m(v_f-v_o)^2\]but since vf=0 we get\[fd=\frac12mv_0^2\]

OpenStudy (anonymous):

I got 0.860 for vo, i forgot the 1/2.

OpenStudy (turingtest):

that's why I'm gonna show you why it's better to keep numbers until the end I also should mention that it helps with sigfigs too....

OpenStudy (turingtest):

so that's\[ W=\Delta K\]\[fd=m\Delta v\]\[\mu N=\frac12m(\cancel{v_f}^0-v_i)^2=\frac12mv_i^2\]\[\mu \cancel mg=\frac12\cancel mv^2\]\[\mu g=v^2\implies v=\sqrt{\mu g}\]so all that using of the mass in our calculations is totally an unnecessary waste of time

OpenStudy (anonymous):

Hey Turing I need to leave for soccer practice, so I'll get back in 2 hours or so to complete this question...sorry :/

OpenStudy (turingtest):

that's cool, while in soccer meditate on the fact that m didn't need to be there and yeah, realize the problem is not done; that's not the velocity we actually want

OpenStudy (turingtest):

typo I forgot the 2 now lol\[ W=\Delta K\]\[fd=m\Delta v\]\[\mu N=\frac12m(\cancel{v_f}^0-v_i)^2=\frac12mv_i^2\]\[\mu \cancel mg=\frac12\cancel mv^2\]\[2\mu g=v^2\implies v=\sqrt{2\mu g}\]

OpenStudy (turingtest):

it magically.... oops :/

OpenStudy (anonymous):

I put it in the wrong place!

OpenStudy (turingtest):

\[W=\Delta K\]\[fd=m\Delta v\]\[\mu Nd=\frac12m(\cancel{v_f}^0-v_i)^2=\frac12mv_i^2\]\[\mu \cancel mgd=\frac12\cancel mv^2\]\[2\mu gd=v^2\implies v=\sqrt{2\mu gd}\]what mistake? I don't know what you mean. (...:P)

OpenStudy (anonymous):

they keep creeping in!

OpenStudy (turingtest):

but now I think we're good from here it is a momentum problem

OpenStudy (anonymous):

Interesting discussion here on this problem!

OpenStudy (turingtest):

indeed, do you know how to proceed from here? if not I will be back in 10 min

OpenStudy (turingtest):

oh you are not the asker lol

OpenStudy (anonymous):

correct.

OpenStudy (turingtest):

well that proves my level of absent-mindedness I'm glad you enjoyed me acting the fool, but even fools must eat later!